A school gym has a perimeter of 324 feet and an area of 6,480 square feet. What are the dimensions of the gym?
Answer by TimothyLamb(4379)

2L + 2w = 324
L = (324 - 2w)/2
Lw = 6480
---
Lw = 6480
w(324 - 2w)/2 = 6480
324w - 2ww = 6480*2
2ww - 324w + 6480*2 = 0
ww - 162w + 6480 = 0
---
the above quadratic equation is in standard form, with a=1, b=-162 and c=6480
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -162 6480
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 90
w = 72
---
either root will work for this problem:
---
w = 72
L = (324 - 2w)/2
L = (324 - 2*72)/2
L = 90
---
answer:
w = 72 ft
L = 90 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php